# Debt, Growth, and the Carbon Lock-In

### v1

@@ B.3 Comparing path dependent intensities to yearly intensities
If we now add stochasticity and use $\pi_{t}=\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right):=\prod_{1\leq\tau^{\prime}\leq\tau}V_{\tau^{\prime}}$, we have $C_{\tau}=LC_{\text{ini}}\pi(\tau)$. In addition, with $X_{\tau^{\prime}}=\ln(V_{\tau^{\prime}})$, we have that $C_{\tau}=LC_{\text{ini}}\prod_{1\leq\tau^{\prime}\leq\tau}e^{X_{\tau^{\prime}}}$. In the limit where $X_{\tau^{\prime}}$ is drawn according to a normal distribution with mean $r$ and standard deviation $\sigma_{X}$, we write $X_{\tau^{\prime}}=r+\sigma_{X}\xi_{\tau^{\prime}}$ ($\xi_{\tau^{\prime}}$ is normally distributed with mean 0 and standard deviation $1$). We can thus write
@@
$\begin{split}\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)&=\frac{\sum_{\tau^{\prime}=0}^{\tau}\left(I(\tau^{\prime})-I(\tau)\right)e^{r\tau^{\prime}}\prod_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}e^{\sigma_{X}\xi_{\tau^{\prime\prime}}}}{\sum_{\tau^{\prime}=0}^{\tau}e^{r\tau^{\prime}}\prod_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}e^{\sigma_{X}\xi_{\tau^{\prime\prime}}}}\\ &=\frac{\sum_{\tau^{\prime}=0}^{\tau}\left(I(\tau^{\prime})-I(\tau)\right)\exp\left(r\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)}{\sum_{\tau^{\prime}=0}^{\tau}\exp\left(r\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)}.\end{split}$ (56)
@@
Since $\xi(\tau)$ is normally distributed, we get that $\left<\exp\left(r\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)\right>=\exp\left((r+\sigma_{X}^{2}/2)\tau^{\prime}\right)$, and
$\mathbb{V}(\exp\left(r\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right))=\exp\left((2r+\sigma_{X}^{2})\tau^{\prime}\right)\left(\exp\left(\sigma_{X}^{2}\tau^{\prime}\right)-1\right).$ (57)
@@
Therefore this variable has wide fluctuations. For small $\sigma_{X}$, we recover the previous deterministic result. We show the dependency of this on the fluctuations $\sigma_{X}$ on Fig.9B. We observe that the difference between path dependent intensity and yearly intensity increases with fluctuations $\sigma_{X}$. This results from the fact that if a phase of large economic growth with large carbon intensities is followed by a phase of recessions with lower carbon intensities (years with a stochastically low growth rate), the weight of carbon-emissive years will be larger. Therefore, the weights of years with large carbon intensities will be larger and the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ will increase, leading to a larger decoupling. In addition, large fluctuations may lead to years with a stochastically large growth rate and low carbon intensities, typically reducing the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$. However, as shown in the deterministic case, the difference typically goes as $1/r$, so that decreasing the growth rate stochastically has a larger impact on decoupling than increasing the growth rate stochastically does. Therefore, recessions lead to an increased decoupling between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ and the average decoupling is larger when fluctuations are larger.
@@
To better understand this effect we can study what happens for a first phase with a high growth rate $\gamma_{h}$, followed by a phase of recession with a lower growth rate $\gamma_{l}$. In this case, for a phase of fast growth lasting for $\tau_{h}$ and a phase of recession lasting for $\tau-\tau_{h}$, we get
@@
$\begin{split}\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)&=\frac{(e^{\gamma_{h}}-1)(e^{\gamma_{l}}-1)}{(e^{\gamma_{l}}-1)(e^{\gamma_{l}(\tau_{h}+1)}-1)+(e^{\gamma_{h}}-1)(e^{\gamma_{h}(\tau+1)}-e^{\gamma_{h}(\tau_{h}+1)})}\times\\ &\left(\frac{\left((e^{\gamma_{h}}-1)((\tau-\tau_{l})e^{\gamma_{h}(\tau_{h}+1)}-\tau)-e^{\gamma_{h}}(1+e^{\gamma_{h}\tau_{h}})\right)}{(e^{\gamma_{h}}-1)^{2}}+\frac{\left((e^{\gamma_{l}}-1)(\tau_{l}-\tau)e^{\gamma_{l}(\tau_{h}+1)}-e^{\gamma_{l}(\tau+1)}+e^{\gamma_{l}(\tau_{h}+1)}\right)}{(e^{\gamma_{l}}-1)^{2}}\right),\end{split}$ (58)
So that for large $\tau$, it scales as the duration of the recession divided by $e^{\lambda_{l}}-1$
@@
Therefore, only the growth rate and duration of the recession phase matter to explain the decoupling between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ in this limit.
@@ B.4 Study of the solvency probability
We can eventually relate this to the probability that a biased random walk is smaller than a threshold at payback time $t_{p}$ [52, 25] (the jumps do not have a zero mean, and the random walk is thus biased). We emphasize that bankruptcy does not correspond to the existence of a time where the debt is larger than the gross capital (or first passage time), but to the fact that at the time of payback $t_{p}$, the debt is larger than the gross capital.
By defining $C_{\tau^{\prime}}=\ln(\gamma_{\tau^{\prime}}/\rho)$ and $S_{\tau}=\sum_{\tau^{\prime}\leq\tau}C_{\tau^{\prime}}$. We assume that the asymetrical distribution of $X$ is $\phi$, with finite mean $m=W-\ln(\rho)$ and variance $\sigma_{X}^{2}$. Using the central limit theorem for large enough $\tau$, we finally obtain that
@@
$\begin{split}P_{S}^{A}(t_{p})&=\mathbb{P}\left(\sum_{\tau^{\prime}\leq\tau}C_{\tau^{\prime}}\geq\ln\left(1-\frac{1}{L}\right)\right)\\ &=\frac{1}{2}\left(1+\text{erf}\left(\frac{t_{p}m-\ln\left(1-\frac{1}{L}\right)}{\sqrt{2t_{p}}\sigma_{X}}\right)\right)\end{split}$ (60)
We can study the argument of the error function and how it depends on time. In particular, if $m>0$ ( $W>\ln(\rho)$), it is a non-monotonic function, decreasing until $t_{p}=-\ln(1-1/L)/(W-\ln(\rho))$ (worst payback time) and then increasing towards $1$. Otherwise, it is a strictly decreasing function of time. Therefore, the result will strongly depend on the sign of $m$.
@@
Figure 10: Solvency probability with Gaussian jumps, but the same average and standard deviation as in Kelly’s case in the limit of large leverage $L=10$ with strategy A.
@@
We plot the solvency probability as a function of the payback time $t_{p}$ in Fig.10 in the limit of large leverage. The short-term growth rate $w(t_{p})$ for strategy A is shown in Fig.11. Using the expression of the short-term growth rate with strategy A, we have that this is defined for all times if $E\geq\rho$. Otherwise this is well defined for all times lower than $\ln(1-1/L)/(\ln(\rho)-\ln(E))$. In addition, if we define
$g(t_{p})=E^{t_{p}}-(1-\frac{1}{L})\rho^{t_{p}},$ (61)
which is the argument of the $\ln$ in the definition of the short-term growth rate, we have that
$w^{\prime}(t_{p})=\frac{1}{t_{p}^{2}}\left(t_{p}\frac{g^{\prime}(t_{p})}{g(t_{p})}-\ln(g(t_{p}))\right).$ (62)
@@
The latter is equal to 0 if and only if $t_{p}g^{\prime}(t_{p})/g(t_{p})=\ln(g(t_{p}))$. Concerning the solvency probability, we observe that this is indeed a decreasing function of the leverage (higher leverage is associated with higher risk). We thus identify four regimes based on the concavity inequality $\ln(E)\geq W$ :
@@
  • If $m>0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 1$, faster as $\sigma_{X}$ decreases. The solvency probability is a non monotonic function of $t_{p}$, first decreasing until $t_{p}=-\ln(1-1/L)/(W-\ln(\rho))$, then increasing towards $1$. For large leverage the first regime vanishes (when $\ln(1-1/L)/(W-\ln(\rho))<1$), but it means that for small enough leverage it may be safer to pay back as soon as possible, or to wait for long enough. The short-term growth rate is well defined at all times.
@@
  • If $m=0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)=\frac{1}{2}$. In this case the solvency probability is decreasing with $t_{p}$. The short term growth rate is well defined at all times.
@@
  • If $m<0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 0$, faster as $\sigma_{X}$ decreases. The solvency probability is decreasing with $t_{p}$. The short term growth rate is well defined at all times.
@@
  • If $m<0$ and $\ln(E)<\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 0$, faster as $\sigma_{X}$ decreases. The solvency probability is decreasing with $t_{p}$. The short term growth rate is only defined for $t_{p}<-\ln(1-1/L)/(\ln(\rho)-\ln(E))$.
Those results are summed up on Fig.11.
Figure 11: solvency probability and short-term growth rate for strategy A, different values of leverage and interest rates. For the solvency probability, there is a transition at $\ln(\rho)=W$. When $W<\ln(\rho)$, the solvency probability is strictly decreasing for all times; otherwise, it is a non-monotonic function of time, with a minimum at $t_{p}=-\ln(1-1/L)/(W-\ln(\rho))$. For the short-term growth rate, there is a transition at $\rho=E$. When the interest rate increases above $E$ and remains below $E^{L/(L-1)}$, the short term growth rate starts decreasing with time after $t_{p}>\ln((1-1/L)\ln(\rho)/\ln(E))/\ln(E/\rho)$. The figure on the left corresponds to $\ln(\rho)<W$. The figure on the right corresponds to $W<\ln(\rho)$
@@
Using strategy B, the solvency probability is slightly different, and we can’t used the central limit theorem anymore
$\begin{split}P_{S}^{B}(t_{p})&=\mathbb{P}\left(\prod_{1\leq\tau^{\prime}\leq t_{p}}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)-\sum_{0\leq t\leq t_{p}-1}\frac{L-1}{\rho}\prod_{1\leq\tau^{\prime}\leq t}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\geq 1-\frac{1}{L}\right).\end{split}$ (63)
This suggests to introduce $\Pi_{t}=\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right)/\rho$, so that
@@
and
@@
If we define the quantity (in terms of the capital $C_{t_{p}}$ and debt $D_{t_{p}}$ at time $t_{p}$)
@@
Therefore, using that the $V_{\tau}$ and $\gamma_{\tau}$ are independent, we get that in average the sign of $X_{t_{p}+1}-X_{t_{p}}$ is set by the sign of $\left<\gamma\right>-\rho$. We can thus draw two conclusions
  • If $\left<\gamma\right><\rho$ (which seems to be the case in most countries), we have that if the debt is larger than the capital at time $t_{p}$ then it is also the case at time $t_{p}+1$ in average. For a fully deterministic model it would mean that the survival probability is a decreasing function of leverage. However due to stochasticity, this result is no longer true when $\gamma$ has a broad distribution (the distribution is broader for higher $\lambda$).
@@
  • Else if $\left<\gamma\right>>\rho$, we have that if the debt is lower than the capital at time $t_{p}$ then it is also the case at time $t_{p}+1$ in average
@@
We plot it on Fig.12, where it reaches a plateau for small enough interest rates. We see that solvency is smaller with strategy B, a risk that has to be taken to sustain a larger growth rate. For $L\sim 1$ we understand why this solvency probability (strategy B) is smaller than for strategy A, as the lower bound for solvency is larger.

### v2

@@ B.3 Comparing path dependent intensities to yearly intensities
Let us now add stochasticity and introduce the stochastic variable $V_{\tau}=\gamma_{\tau}+L-1$ and $\pi_{t}=\prod_{1\leq\tau^{\prime}\leq\tau}V_{\tau^{\prime}}$ so that the capital is $C_{\tau}=LC_{\text{ini}}\pi(\tau)$. Let us also introduce the random variable $X_{\tau^{\prime}}=\ln(V_{\tau^{\prime}})$, so that $C_{\tau}=LC_{\text{ini}}\prod_{1\leq\tau^{\prime}\leq\tau}e^{X_{\tau^{\prime}}}$. In the limit where $X_{\tau^{\prime}}$ is drawn according to a normal distribution with mean $W$ and standard deviation $\sigma_{X}$, we write $X_{\tau^{\prime}}=W+\sigma_{X}\xi_{\tau^{\prime}}$ ($\xi_{\tau^{\prime}}$ is normally distributed with mean 0 and standard deviation $1$). We can thus write
@@
$\begin{split}\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)-I(\tau)&=\frac{\sum_{\tau^{\prime}=0}^{\tau}\left(I(\tau^{\prime})-I(\tau)\right)e^{W\tau^{\prime}}\prod_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}e^{\sigma_{X}\xi_{\tau^{\prime\prime}}}}{\sum_{\tau^{\prime}=0}^{\tau}e^{W\tau^{\prime}}\prod_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}e^{\sigma_{X}\xi_{\tau^{\prime\prime}}}}\\ &=\frac{\sum_{\tau^{\prime}=0}^{\tau}\left(I(\tau^{\prime})-I(\tau)\right)\exp\left(W\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)}{\sum_{\tau^{\prime}=0}^{\tau}\exp\left(W\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)}.\end{split}$ (58)
@@
Since $\xi(\tau)$ is normally distributed, we get that $\left<\exp\left(W\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)\right>=\exp\left((W+\sigma_{X}^{2}/2)\tau^{\prime}\right)$, and
$\mathbb{V}\left(\exp\left(W\tau^{\prime}+\sigma_{X}\sum_{1\leq\tau^{\prime\prime}\leq\tau^{\prime}}\xi_{\tau^{\prime\prime}}\right)\right)=\exp\left((2W+\sigma_{X}^{2})\tau^{\prime}\right)\left(\exp\left(\sigma_{X}^{2}\tau^{\prime}\right)-1\right).$ (59)
@@
Therefore, this variable has wide fluctuations. For small $\sigma_{X}$, we recover the previous deterministic result. We show the dependency of this on the fluctuations $\sigma_{X}$ in Fig.9B. We observe that the difference between the path-dependent intensity and the yearly intensity increases with the fluctuations $\sigma_{X}$. This results from the fact that if a period of high economic growth with high carbon intensities is followed by a period of recessions with lower carbon intensities (years with a stochastically low growth rate), the weight of carbon-emissive years will be larger. Therefore, the weights of years with large carbon intensities will be larger and the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ will increase. In addition, large fluctuations may lead to years with a stochastically significant growth rate and low carbon intensities, typically reducing the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$. However, as shown in the deterministic case, the difference typically scales as $1/W$, so that decreasing the growth rate stochastically has a larger impact on the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ than increasing it stochastically. Therefore, recessions lead to a enhanced difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$, which average rises when fluctuations are stronger.
@@
To better understand this effect, we can study what happens for a first phase with a high growth rate $\gamma_{h}$, followed by a phase of recession with a lower growth rate $\gamma_{l}$. In this case, for a phase of fast growth lasting for $\tau_{h}$ and a phase of recession lasting for $\tau-\tau_{h}$, we get
@@
$\begin{split}\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)&-I(\tau)=\frac{(e^{\gamma_{h}}-1)(e^{\gamma_{l}}-1)}{(e^{\gamma_{l}}-1)(e^{\gamma_{l}(\tau_{h}+1)}-1)+(e^{\gamma_{h}}-1)(e^{\gamma_{h}(\tau+1)}-e^{\gamma_{h}(\tau_{h}+1)})}\times\\ &\left(\frac{\left((e^{\gamma_{h}}-1)((\tau-\tau_{l})e^{\gamma_{h}(\tau_{h}+1)}-\tau)-e^{\gamma_{h}}(1+e^{\gamma_{h}\tau_{h}})\right)}{(e^{\gamma_{h}}-1)^{2}}+\frac{\left((e^{\gamma_{l}}-1)(\tau_{l}-\tau)e^{\gamma_{l}(\tau_{h}+1)}-e^{\gamma_{l}(\tau+1)}+e^{\gamma_{l}(\tau_{h}+1)}\right)}{(e^{\gamma_{l}}-1)^{2}}\right),\end{split}$ (60)
So that for large $\tau$, it scales as the duration of the recession divided by $e^{\gamma_{l}}-1$
@@
Therefore, only the growth rate and duration of the recession phase matter to explain the difference in trends between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ in this limit.
@@ B.4 Study of the solvency probability
We can eventually relate this to the probability that a biased random walk is below a threshold at payback time $t_{p}$ [57, 30] (the jumps do not have zero mean, and the random walk is thus biased). We emphasize that bankruptcy does not correspond to the existence of a time where the debt is larger than the gross capital (or first passage time), but to the fact that at the time of payback $t_{p}$, the debt is larger than the gross capital.
To evaluate it, let us start by introducing the stochastic variable $X_{\tau^{\prime}}=\ln(\gamma_{\tau^{\prime}}/\rho)$, which has mean $m=W-\ln(\rho)$ and variance $\sigma_{X}^{2}$. Using the central limit theorem for large enough $\tau$, the solvency probability is
@@
$\begin{split}P_{S}^{A}(t_{p})&=\mathbb{P}\left(\sum_{\tau^{\prime}\leq\tau}X_{\tau^{\prime}}\geq\ln\left(1-\frac{1}{L}\right)\right)\\ &=\frac{1}{2}\left(1+\text{erf}\left(\frac{t_{p}m-\ln\left(1-\frac{1}{L}\right)}{\sqrt{2t_{p}}\sigma_{X}}\right)\right).\end{split}$ (62)
@@
Figure 10: Solvency probability with Gaussian jumps, but the same average and standard deviation as in Kelly’s case, in the limit of large leverage $L=10$ with strategy A.
@@
From this expression, we find that the solvency probability strongly depends on the sign of $m$. We plot this solvency probability as a function of the payback time $t_{p}$ in Fig.10 in the limit of large leverage. Using the expression of the short-term growth rate with strategy A given in Eq. 19, we have that this is defined for all times if $E\geq\rho$. Otherwise, this is well defined for all times lower than $\ln(1-1/L)/(\ln(\rho)-\ln(E))$.
@@
Regarding the solvency probability, we observe that it is a decreasing function of leverage (higher leverage is associated with higher risk). We thus identify four regimes based on the concavity inequality $\ln(E)\geq W$ :
@@
  • If $m>0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 1$, faster as $\sigma_{X}$ decreases. The solvency probability is a non monotonic function of $t_{p}$, first decreasing until $t_{p}=-\ln(1-1/L)/(W-\ln(\rho))$, then increasing towards $1$. For large leverage, the first regime vanishes (when $\ln(1-1/L)/(W-\ln(\rho))<1$), but it means that for small enough leverage, it may be safer to pay back as soon as possible, or to wait for long enough. The short-term growth rate is well defined at all times.
@@
  • If $m=0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)=\frac{1}{2}$. In this case, the solvency probability is decreasing with $t_{p}$. The short-term growth rate is well defined at all times.
@@
  • If $m<0$ and $\ln(E)\geq\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 0$, faster as $\sigma_{X}$ decreases. The solvency probability is decreasing with $t_{p}$. The short-term growth rate is well defined at all times.
@@
  • If $m<0$ and $\ln(E)<\ln(\rho)$, $P_{S}^{A}(t_{p}\to\infty)\to 0$, faster as $\sigma_{X}$ decreases. The probability of solving decreases with $t_{p}$. The short-term growth rate is only defined for $t_{p}<-\ln(1-1/L)/(\ln(\rho)-\ln(E))$.
@@
Using strategy B, the solvency probability has a more complex expression, and we can not simply it directly with the central limit theorem. From Eq. 32, we have:
$\begin{split}P_{S}^{B}(t_{p})&=\mathbb{P}\left(\prod_{1\leq\tau^{\prime}\leq t_{p}}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)-\sum_{0\leq t\leq t_{p}-1}\frac{L-1}{\rho}\prod_{1\leq\tau^{\prime}\leq t}\frac{1}{\rho}\left(\gamma_{\tau^{\prime}}+(L-1)\right)\geq 1-\frac{1}{L}\right),\end{split}$ (63)
which suggests to introduce the quantity $\Pi_{t}=\prod_{1\leq\tau^{\prime}\leq t}\left(\gamma_{\tau^{\prime}}+(L-1)\right)/\rho$, so that
@@
which is Eq. (14) of the main text with $X_{t_{p}}=\Pi_{t_{p}}-\frac{L-1}{\rho}\sum_{0\leq t\leq t_{p}-1}\Pi_{t}$.
Then, we have:
@@
Now, we can rewrite $X_{t_{p}}$ in terms of the capital $C_{t_{p}}$ and debt $D_{t_{p}}$ at time $t_{p}$ as
@@
Therefore, using the fact that the $V_{\tau}$ and $\gamma_{\tau}$ are independent, we get that on average the sign of $X_{t_{p}+1}-X_{t_{p}}$ is set by the sign of $\left<\gamma\right>-\rho$. We can thus draw two conclusions
  • If $\left<\gamma\right><\rho$ (which seems to be the case in most countries), we have that if the debt is larger than the capital at time $t_{p}$, then it is also the case at time $t_{p}+1$ on average. For a fully deterministic model, it would mean that the survival probability is a decreasing function of leverage. However, due to stochasticity, this result is no longer true when $\gamma$ has a broad distribution (the distribution is broader for higher $W$).
@@
  • Otherwise if $\left<\gamma\right>>\rho$, we have that if the debt is lower than the capital at time $t_{p}$ then it is also the case at time $t_{p}+1$ in average
@@
We plot it on Fig.11, where it reaches a plateau for small enough interest rates. We see that solvency is smaller with strategy B, a risk that has to be taken to sustain a larger growth rate. For $L\sim 1$ we understand why this solvency probability (strategy B) is smaller than for strategy A, because the investor in strategy B relies more on leverage than in strategy A.